OpenStudy (anonymous):

2x-3y=5 10x+10y=30 what is the solution?

OpenStudy (anonymous):

Multiply first equation by 5. Then add.

myininaya (myininaya):

how about -5 and then after add the equations

myininaya (myininaya):

5 is fine we would just have to subtract which is fine

OpenStudy (anonymous):

It should be subtract. Anyway, once you multiply by 5, it will be clear whether to add or subtract to eliminate x. At least I hope.

OpenStudy (anonymous):

please elaborate

myininaya (myininaya):

\[-10x+30y=-25\] multiplying the top equation by -5 gives us that

myininaya (myininaya):

now add this to 10x+10y=30

OpenStudy (anonymous):

whats the answer

myininaya (myininaya):

add those equations (-10x+10x)+(30y+10y)=(-25+30)

myininaya (myininaya):

\[40y=5 => y=\frac{5}{40}\] reduced then plug into one of the equations given and solve for x

myininaya (myininaya):

sorry i multiplied wrong at first the equation was suppose to be -10x+15y=-25 after multiplying the -5

myininaya (myininaya):

now add the equations together

OpenStudy (anonymous):

i dont remember my teacher explaining it like that.. o.o

myininaya (myininaya):

\[-10x+15y=-25\] \[10x+10y=30\] ---------------- 0x+25y=5 25y=5 now divide by 25 on both sides then reduce the fraction then plug into one of the equations given and solve for x